# Can Large Language Models Unveil the Mysteries? An Exploration of Their Ability to Unlock Information in Complex Scenarios

### v1

@@ 1 Introduction
Prompt engineering has recently become the popular method for enhancing model performance, particularly in Chain of Thought (COT) reasoning [49] and reflection prompt [31]. To explore its role in combinatorial reasoning tasks that integrate multiple perceptual inputs in complex scenarios, we conduct tests by manual prompts to stimulate more thoughtful model responses and select 36 complex scenarios from the game ‘Can you escape?’ [38], such as “combination of reasoning between color sequences and colored numbers” in Figure 1, which require combinatorial reasoning across multiple perceptual information. To encourage thoughtful responses, we utilize manual prompts to guide the model’s most confident choices, helping it gradually integrate multiple perceptual inputs and complete combinatorial reasoning, as shown in Figure 2. Specifically, the model determines whether to continue exploring the scene or proceed with a final action through five binary classification tasks, each featuring distinct prompt designs. A joint calibration set is constructed based on the original calibration set and the confidence interval from conformal prediction [56], enabling the calculation of inconsistency degrees. The action corresponding to the minimum inconsistency is then selected as the next step. If the model opts to continue exploring, an artificial cue is introduced, and the scene is described in detail. Alternatively, if the model decides to resolve the mystery, all scene images and cues are utilized for prediction. Experimental results show that, despite prompt engineering for ChatGPT-4o [24], reasoning failed in 36 scenarios, underscoring prompt engineering’s limited impact and details are in Sec. 5.1.
@@
To address the absence of evaluation of combining multiple perceptual inputs and performing combinatorial reasoning in complex scenarios, we propose two benchmarks: 1) Clue-Visual Question Answering (CVQA) comprises three task types designed to test a model’s ability to integrate textual and visual perceptual information to identify relevant clues from unspecified visual data. The first task involves reasoning with a combination of unspecified visual entities located in different scenes that can interact, such as a “key” in one scene that can be used to unlock a “lock” in another scene. The second task involves reasoning with specified textual entities and unspecified visual entities, where the entities mentioned in the text can interact with a certain combination of visual entities in complex scenarios. The third task utilizes specified sequence types (e.g., number sequences, letter sequences, etc.) to combine all visual scenes to infer clues, such as ‘Four colored numbers’ and ‘Four color sequences’ in Figure 1. 2) Clue of Password-Visual Question Answering (CPVQA) consists of two task types that evaluate the model’s ability to combine visual perceptual information for accurate combinatorial reasoning. These tasks involve reasoning across all visual scenes to determine the final numeric sequence (e.g., 2536 in Figure 2) or entity code sequence (e.g., BCDA in Figure 3). In particular, our benchmarks include several unique features: 1) The questions in our benchmarks are designed to guide the model to achieve combinatorial reasoning in complex scenarios. 2) The topics in each of our complex scenarios remain consistent by selecting visually similar images from the same game level [38] rather than from existing datasets. 3) Our benchmarks’ scale is manually expanded to produce high-quality evaluations, with the effectiveness of this approach demonstrated in Sec. 3.2.
We propose three methods to bypass prompt engineering, aiming to explore and enhance the model’s ability to integrate multiple perceptual inputs for combinatorial reasoning in complex scenarios. Most current MLLMs effectively process only single-image inputs, so this study assumes the model accepts one image as input: 1) Contextual learning [32] has proven effective. To ensure the model processes all images during reasoning, MLLMs generate objective, detailed captions as rich sources of image information. These captions, along with questions, are input into LLMs and MLLMs for reasoning. A main scene image is selected to represent the overall theme of the complex scenario. 2) To avoid the labor of designing unique COT prompts for each scenario, the probability difference in predicted logits is leveraged to enhance reasoning [47]. Utilizing a minimum margin decoding strategy, confidence is assessed by comparing the top two options, then high-confidence tokens are introduced to generate coherent answers. 3) Building on recent advancements [16, 18], we propose an image-based semantic retrieval method to enhance visual detail for MLLMs. An advanced LLM predicts the relevance between questions and image captions, with the most relevant caption guiding scene selection for MLLM inference.
@@
We make a thorough evaluation of various well-known MLLMs and LLMs on our proposed benchmarks. In CVQA, results of model inference (Sec. 4.1) are suboptimal, notably, COT reasoning without prompting (Sec. 4.2) and semantic retrieval (Sec. 4.3) improve performance by up to 19%. In CPVQA, performance falls short of expectations. Overall, while our method enhances combinatorial reasoning in complex scenarios, model accuracy remains low. Our contributions can be summarized as follows: 1) A comprehensive examination reveals that even SOTA models like ChatGPT-4o perform suboptimally in integrating multiple perceptual inputs for combinatorial reasoning in complex scenarios. 2) To our knowledge, this is the first study for combinatorial reasoning in complex scenarios requiring multiple perceptual inputs, with methods designed to bypass prompt engineering. 3) The proposed plug-and-play method achieves comparable improvements over SOTA closed-source models.
@@ 3.1 Benchmark Structure
Image source. In this study, we selected the game ‘Can You Escape?’ [38] as our image source due to its close resemblance to real-world visuals. The image named “scene1” in each image set is the main scene, and the other images are details of this scene. In addition to visual realism, our benchmarks offer several significant advantages: 1). Unified scene themes: While most existing datasets treat a single image as an independent scene, our benchmarks utilize multiple images with the same theme to create a cohesive and complete scene. 2). Diverse task types: Our benchmarks encompass a variety of reasoning tasks within the same scenario, including reasoning about relationships between entities, interpreting digital codes, and rearranging sequences based on colors and shapes. 3). Strong adaptability: We utilize advanced image processing tools to modify and augment images, a process that would be costly in real-world scenarios. This approach allows us to expand our benchmarks in a cost-effective manner.
@@
  • The password task requires the model to infer a string of letters and numbers from a scene composed of multiple images. The question specifies the components (letters, numbers or a combination) and the number of digits in the final result. In some images, character combinations may appear in different forms, such as Arabic numerals or Roman numerals. It is important to note that in certain images, the final character sequence may need to be reorganized based on attributes like color, shape, or other distinguishing features.
  • The sequence rearrangement task involves determining the correct order of entities in a given image (e.g., the order of the trophies in Figure 3) within a scene composed of multiple images. The entities are encoded based on their order in the image from the question (e.g., ‘ABCD’ from left to right). However, the final encoded sequence is influenced by information from other images, such as color, shape, or other attributes. To account for the possibility that the language model may repeat the question, we have designed the final true sequence to differ from the initial encoded sequence designed in the question.
3.2 Extension Method
@@ 4.3 Semantic Retrieval
In this subsection, we introduce a semantic retrieval method for handling multiple image inputs. While the main scene in model inference (Sec. 4.1) of this study provides the MLLM with key information, such as the theme of the scene as the image input, it may not always be the image most relevant to the answer. Therefore, across all known scenes, we evaluate the semantic relevance between the question and the scene descriptions to identify the image most likely to be relevant to the answer. This study implements two methods:
@@ 5 Experiments
This section first introduces the details of exploratory experiment (Sec. 5.1), which utilizing the high-level reasoning model to perform combinatorial reasoning in complex scenarios. Next, we describe the experimental setup (Sec. 5.2), which includes the benchmark evaluation mechanism and the models used in the experiments. Finally the main results (Sec. 5.3) of our proposed methods are presented.
@@ 5.1 Exploratory Experiment
In this subsection, we describe the process of constructing artificial prompts to guide the model through combinatorial reasoning step by step. The specific steps are as follows:
@@ 5.2 Experimental Setup
Models. In this paper, we utilize four open-source and six closed-source models to evaluate the performance of various models and configurations. 1). For caption generation, we utilize three open-source models: LLaVA-1.6-7b [19], LLaVA-1.6-13b [19], and Qwen-VL [44]. 2). For model reasoning tasks, we utilize open-source models such as Mistral-7b [13], LLaVA-1.6-7b, LLaVA-1.6-13b, and Qwen-VL. Additionally, we utilize closed-source models including ChatGPT-3.5-turbo [24], ChatGPT-4, ChatGPT-4o, ChatGPT4V, Gemini Pro [36], and Qwen-Max [3].
@@ 5.3 Main Results
We first present the model score results for method model inference (Sec. 4.1) in Tables 1 and 2, and then present the score results of our improved methods—COT reasoning (Sec. 4.2) and semantic retrieval (Sec. 4.3)—through ablation experiments shown in Tables 3 and 4, respectively.
@@
Table 1: The main results of method LLMs reasoning (Sec. 4.1). In the tables of this paper, i: The bold font is the first place in the same task category, and the underlined font is the second place; ii: ‘Caption’ indicates the type of model that generates captions; iii: And ‘all’ represents the total score of all tasks in the benchmark, ‘pc’ represents prop collection , ‘pu’ represents props usage, ‘pcl’ represents password clues, ‘pas’ represents password, and ‘seq’ represents sequence rearrangement.
@@
Ablation on COT reasoning without prompting. In Table 3, we utilize the ablation method to demonstrate the effect of COT reasoning without prompting in method COT reasoning without prompting (Sec. 4.2). 1) In the LLM reasoning method, the open-source model Mistral-7b was utilized as the base model. Table 3 presents the COT reasoning method significantly improved the model’s performance on the CVQA benchmark. However, it only partially enhance performance on the CPVQA benchmark. 2) In the MLLM reasoning method, this paper explores models with different parameter sizes of LLaVA-1.6. We find that in CVQA, the COT reasoning method had little to no effect on improving the performance of MLLMs. In CPVQA, COT reasoning method led to greater improvements for models with larger parameters (13b), which may be related to the random design of our COT reasoning method. Since the number of parameters is proportional to the vocabulary size, models with more parameters have a higher likelihood of generating a complex COT reasoning path. The COT reasoning method may be more effective for text reasoning and models with larger parameters, likely due to its inherent randomness.
@@
Table 4: The ablation results of method semantic retrieval (Sec. 4.3), where the DOI indicates Description of Other Images. The ‘DOI’ in the table, i(✓) combines other image captions in the final prediction; ii(✗) excludes combination, aligning with semantic retrieval (Sec. 4.3); iii(-)uses only MLLM for reasoning.
Ablation on semantic retrieval. We compare the results of three methods in Table 4: I: Results of method MLLM reasoning (Sec. 4.1). II: The method utilizing inputs $\mathcal{C}$ and $v^{\prime}$, and III: the method utilizing only $v^{\prime}$ as input, without $\mathcal{C}$, in Sec. 4.3. Specifically, in CVQA, method III shows a clear advantage, particularly in the case of Model: Qwen-VL  Caption: Qwen-VL, where its scores are consistently optimal. In contrast, method II yields suboptimal results. This indicates that in CVQA, our proposed method in Sec. 4.3 offers an improvement over the direct use of MLLM reasoning. In CPVQA, method III performs slightly better than the other methods, while method I yields suboptimal results. Our propose method offers an improvement over direct model prediction; however, more refined methods are needed to further enhance performance.
Additional Experiments. Given that models capable of handling multi-image inputs exist, the computational resources required to locally deploy open-source models increase significantly as the number of images grows. To maintain the comprehensiveness of our experiment, we utilize API-callable multi-image input models to directly evaluate their performance on our proposed benchmarks. The results are presented in Table 5.
@@
In-depth analysis. In our exploration and experiments, the primary error observed is that it tends to answer based solely on a single information source, without integrating other relevant data for joint reasoning, as shown in Figure 2. Additionally, for certain questions, such as the trophy sequence in Figure 3, the model tends to respond based only on the information presented in the question. This highlights the potential to further integrate multiple information for combined reasoning. Furthermore, as shown in Tables 3 and 5, models utilizing COT or trained on COT data demonstrate better performance. This suggests that enhanced reasoning methods, such as COT [47, 49], TOT [53], and RL-trained models [6], may be more effective in analyzing information from multiple sources.
@@ 6 Conclusion
In this work, we introduce two benchmarks, CVQA and CPVQA, designed to evaluate the ability of models to perform combinatorial reasoning by integrating multiple perceptual information in complex contexts, and we propose three baselines methods. Extensive experiments reveal that current models struggle with combinatorial reasoning in complex contexts. Our method alleviates these limitations, improves performance, and provides insights for future model enhancements.
@@ Appendix A Benchmark Solution Formats
In this Section, in addition to the solved forms of CVQA and CPVQA shown in Figure 3, we additionally show three other tasks of CVQA and CPVQA to emphasize the characteristics of different tasks on different benchmarks. This paper specifically quotes these images and explains them in text in Sec 2.1 and Sec 2.2.

### v2

@@ 1 Introduction
Unfortunately, there are currently no datasets or simulation scenarios that combine multiple perceptual information for combinatorial reasoning. However, we find that human often pay attention to different clues and then infer the answer when playing a game called escape room. We try to utilize ChatGPT-4o [24], a powerful closed-source model that can perform reasoning, to conduct exploratory experiments in the game Can you escape?  [38], as shown in Sec. 5.1, and the results failed in reasoning in 36 complex scenarios. To address the lack of evaluation in integrating multiple perceptual inputs and performing combinatorial reasoning in complex scenarios—where the intricacy and accuracy of the task make errors significantly impact the reasoning process—we meticulously curate two benchmarks with enhanced rigor. Specifically, we manually select 155 complex scenarios from this game, encompassing environments commonly encountered in real life (such as schools, laboratories, and stadiums). Furthermore, we systematically modify the pigment-related information in the images, thereby expanding the benchmark size to 2.4 times its original scale.
@@
Specifically, we propose two benchmarks: 1) Clue-Visual Question Answering (CVQA) comprises three task types designed to test a model’s ability to integrate textual and visual perceptual information to identify relevant clues from unspecified visual data. The first task involves reasoning with a combination of unspecified visual entities located in different scenes that can interact, such as a “key” in one scene that can be used to unlock a “lock” in another scene. The second task involves reasoning with specified textual entities and unspecified visual entities, where the entities mentioned in the text can interact with a certain combination of visual entities in complex scenarios. The third task utilizes specified sequence types (e.g., number sequences, letter sequences, etc.) to combine all visual scenes to infer clues, such as ‘Four colored numbers’ and ‘Four color sequences’ in Figure 1. 2) Clue of Password-Visual Question Answering (CPVQA) consists of two task types that evaluate the model’s ability to combine visual perceptual information for accurate combinatorial reasoning. These tasks involve reasoning across all visual scenes to determine the final numeric sequence (e.g., 2536 in Figure 1) or entity code sequence (e.g., BCDA in Figure 2). In particular, our benchmarks include several unique features: 1) The questions in our benchmarks are designed to guide the model to achieve combinatorial reasoning in complex scenarios. 2) The topics in each of our complex scenarios remain consistent by selecting visually similar images from the same game level [38] rather than from existing datasets. 3) Our benchmarks’ scale is manually expanded to produce high-quality evaluations, with the effectiveness of this approach demonstrated in Sec. 3.2.
To enhance the model’s capability in integrating multiple perceptual inputs for combinatorial reasoning in complex scenes, we propose three methodological approaches from distinct perspectives: 1) From the context, contextual learning  [32] has been shown to be effective. We utilize text as a medium for reasoning enhancement in Sec. 4.1. 2) From the model, the probability difference in prediction logic is used to enhance reasoning  [48], as shown in Sec. 4.2. 3) From the data, based on recent progress  [15, 18], we propose an image-based semantic retrieval method in Sec. 4.3 to enhance the visual details of MLLMs.
@@
We make a thorough evaluation of various well-known MLLMs and LLMs on our proposed benchmarks. In CVQA, results of contextual learning (Sec. 4.1) are suboptimal, notably, Chain-Of-Thoughts (COT) reasoning without prompting (Sec. 4.2) and semantic retrieval (Sec. 4.3) improve performance by up to 19%. In CPVQA, performance falls short of expectations. Overall, while our method enhances combinatorial reasoning in complex scenarios, model accuracy remains low. The main problems are the tendency to answer questions based on a single source of information, over-reliance on information in the question, or reasoning based solely on the model’s prior knowledge.
Our contributions can be summarized as follows: 1) To our knowledge, this is the first study for combinatorial reasoning in complex scenarios requiring multiple perceptual inputs. 2) Experimental results indicate that even SOTA closed-source models (such as ChatGPT-4o) fail to complete our proposed benchmark, we analysis of their errors and potential improvements. 3) The proposed plug-and-play method achieves comparable improvements over SOTA closed-source models.
@@ 3.1 Benchmark Structure
Image source. In this study, we selected the game Can You Escape? [38] as our image source due to its close resemblance to real-world visuals. The image named “scene1” in each image set is the main scene, and the other images are details of this scene. In addition to visual realism, our benchmarks offer several significant advantages: 1). Unified scene themes: While most existing datasets treat a single image as an independent scene, our benchmarks utilize multiple images with the same theme to create a cohesive and complete scene. 2). Diverse task types: Our benchmarks encompass a variety of reasoning tasks within the same scenario, including reasoning about relationships between entities, interpreting digital codes, and rearranging sequences based on colors and shapes. 3). Strong adaptability: We utilize advanced image processing tools to modify and augment images, a process that would be costly in real-world scenarios. This approach allows us to expand our benchmarks in a cost-effective manner.
@@
  • Password. This task requires the model to infer a string of letters and numbers from a scene composed of multiple images. The question specifies the components (letters, numbers or a combination) and the number of digits in the final result. In some images, character combinations may appear in different forms, such as Arabic numerals or Roman numerals. It is important to note that in certain images, the final character sequence may need to be reorganized based on attributes like color, shape, or other distinguishing features.
  • Sequence rearrangement. The task involves determining the correct order of entities in a given image (e.g., the order of the trophies in Figure 2) within a scene composed of multiple images. The entities are encoded based on their order in the image from the question (e.g., ‘ABCD’ from left to right). However, the final encoded sequence is influenced by information from other images, such as color, shape, or other attributes. To account for the possibility that the language model may repeat the question, we have designed the final true sequence to differ from the initial encoded sequence designed in the question.
3.2 Benchmark Extension
@@ 4.3 Semantic Retrieval
In this subsection, we introduce a semantic retrieval method for handling multiple image inputs. While the main scene in contextual learning/ (Sec. 4.1) of this study provides the MLLM with key information, such as the theme of the scene as the image input, it may not always be the image most relevant to the answer. Therefore, across all known scenes, we evaluate the semantic relevance between the question and the scene descriptions to identify the image likely to be relevant to the answer. This study implements two methods:
@@ 5 Experiments
This section first introduces the details of exploratory experiment (Sec. 5.1), which utilizing the high-level reasoning model to perform combinatorial reasoning in complex scenarios. Next, we describe the experimental setup (Sec. 5.2), which includes the benchmark evaluation mechanism and the models used in the experiments. Finally the results (Sec. 5.3) and in-depth analysis (Sec. 5.4) of our proposed methods are presented.
@@ 5.1 Exploratory Experiment
We test with manual prompts to stimulate more thoughtful model responses, as shown in Figure  5. The specific steps are as follows:
@@ 5.2 Experimental Setup
Models. In this paper, we utilize four open-source and six closed-source models to evaluate the performance of various models and configurations. 1). For caption generation, we utilize three open-source models: LLaVA-1.6-7b [19], LLaVA-1.6-13b [19], and Qwen-VL [45]. 2). For model reasoning tasks, we utilize open-source models such as Mistral-7b [39], LLaVA-1.6-7b, LLaVA-1.6-13b, and Qwen-VL. Additionally, we utilize closed-source models including ChatGPT-3.5-turbo [24], ChatGPT-4, ChatGPT-4o, ChatGPT-4V, Gemini Pro [36], and Qwen-Max [3].
@@ 5.3 Main Results
We first present the model score results for method contextual learning (Sec. 4.1) in Tables 1 and 2, and then present the score results of our improved methods—COT reasoning (Sec. 4.2) and semantic retrieval (Sec. 4.3)—through ablation experiments shown in Tables 3 and 4, respectively.
@@
Table 1: The main results of method LLMs reasoning (Sec. 4.1). In the tables of this paper, the bold font is the first place in the same task category, and the underlined font is the second place; ‘Caption’ indicates the type of model that generates captions; and ‘all’ represents the total score of all tasks in the benchmark, ‘pc’ represents prop collection , ‘pu’ represents props usage, ‘pcl’ represents password clues, ‘pas’ represents password, and ‘seq’ represents sequence rearrangement.
@@
Ablation on COT reasoning without prompting. In Table 3, we utilize the ablation method to demonstrate the effect of COT reasoning without prompting in method COT reasoning without prompting (Sec. 4.2). 1) In the LLM reasoning method, the open-source model Mistral-7b was utilized as the base model. Table 3 presents the COT reasoning method significantly improved the model’s performance on the CVQA benchmark. However, it only partially enhance performance on the CPVQA benchmark. 2) In the MLLM reasoning method, this paper explores models with different parameter sizes of LLaVA-1.6. We find that in CVQA, the COT reasoning method had little to no effect on improving the performance of MLLMs. In CPVQA, COT reasoning method led to greater improvements for models with larger parameters (13b), which may be related to the random design of our COT reasoning method. Since the number of parameters is proportional to the vocabulary size, models with more parameters have a higher likelihood of generating a complex COT reasoning path. The COT reasoning method may be more effective for text reasoning and models with larger parameters, due to its inherent randomness.
@@
Table 4: The ablation results of method semantic retrieval (Sec. 4.3), where the DOI indicates Description of Other Images. The ‘DOI’ in the table, (✓) combines other image captions in the final prediction; (✗) excludes combination, aligning with semantic retrieval (Sec. 4.3); (-) uses only MLLM for reasoning.
Ablation on semantic retrieval. We compare the results of three methods in Table 4, the MLLM reasoning method in contextual learning (Sec. 4.1), utilizing inputs $\mathcal{C}$ and $v^{\prime}$, and only $v^{\prime}$ as input, without $\mathcal{C}$ (Sec. 4.3). Specifically, in CVQA, the method that relies solely on $v^{\prime}$ as input, without incorporating $\mathcal{C}$, demonstrates a significant advantage. Notably, when utilizing captions generated by Qwen-VL and employing Qwen-VL for reasoning, this approach consistently achieves the highest scores. In contrast, the method utilizing both $\mathcal{C}$ and $v^{\prime}$ yields suboptimal results, suggesting that in CVQA, our approach in semantic retrieval (Sec. 4.3) outperforms contextual learning (Sec. 4.1). Similarly, in CPVQA, the method that relies solely on $v^{\prime}$ without incorporating $\mathcal{C}$ demonstrates a slight performance advantage over other approaches, whereas the MLLM reasoning method within the contextual learning framework produces less than optimal results.
Additional Experiments. Given that models capable of handling multi-image inputs exist. To maintain the comprehensiveness of our experiment, we utilize multi-image input models to directly evaluate their performance on our proposed benchmarks. The results are presented in Table 5.
@@
5.4 In-depth Analysis
In our exploration and experiments, we identify three primary types of errors in the model’s reasoning. First, the model often relies on a single source of information rather than integrating multiple relevant inputs for joint reasoning. For instance, it may generate the answer “6325” without considering the color order depicted in Figure 1, highlighting deficiencies in its multimodal retrieval and integration capabilities. Second, the model tends to over-rely on the explicit content of the query without critically analyzing additional contextual information. For example, in Figure  2, it directly concludes that the final trophy sequence is “ABCD” reflecting a tendency to follow surface-level cues rather than engage in deeper reasoning. Finally, when faced with ambiguous or incomplete information, the model often determines that the provided data is insufficient and instead resorts to prior knowledge for inference. This over-reliance on pre-existing biases underscores its limitations in fine-grained recognition and precise analytical reasoning.
Furthermore, effective methods, as shown in Tables 3 and 5, models using COT or trained on COT data perform better, suggesting that enhanced reasoning methods, such as COT [48, 50] and RL-trained models [6], may be more effective in analyzing information from multiple sources. Another notable observation is that text-based reasoning frequently outperforms joint reasoning with the input image, indicating that the integration of multiple visual information sources warrants further refinement.
@@ 6 Conclusion
We present two benchmarks to assess models’ combinatorial reasoning with multiple cues in complex scenarios. Experiments reveal current models struggle. We propose three baselines that mitigates these limitations, and guides future research. This study reveal key areas for improvement in multimodal reasoning, contextual comprehension, and fine-grained perception, which are crucial for enhancing the model’s robustness in complex scenarios.
Limitations: Due to the dataset’s complexity, we rely on manual annotations to ensure methodological rigor. We anticipate that future research will improve scalability.
Supplementary Material
@@ Appendix A Benchmark Solution Formats
In this Section, in addition to the solved forms of CVQA and CPVQA shown in Figure 2, we additionally show three other tasks of CVQA and CPVQA to emphasize the characteristics of different tasks on different benchmarks. This paper specifically quotes these images and explains them in text in Sec. 2.1 and Sec. 2.2.